In an article about Investment abernatives, Money magazine reported that drig stocks provide a potentisli for long-term growth, longs on a regular basis. For aduts age 65 and oide, 80% take prescrution o the adut population (Stetistical Abstract of the United States, 2008) o Round your answers to 4 decimal places dugs regularly. For owns age is eo 64, 44% take preserwon drugs regularly. The age 18-64 age group accounts for 84.3% of What is the probability that a randomly selected adult is 65 or older .Given an dut takes prescrigtion dngs regularly, what is the probablity that the adut s 6s or older?
Answer #1

65 and older - 80 % for drugs
18-64 - 44 % for drugs
p = 84.3%

a)
P(65 or older ) - 1 - P( 18-64) = 1- 0.843
= 0.157

b)
P(65 or older|take drugs)
= P( drugs and 65-older)/P(take drugs)
= 0.157 * 0.8/(0.157 * 0.8 + 0.843 * 0.44)
= 0.25296

Please rate